# GMAI-VL & GMAI-VL-5.5M: A Large Vision-Language Model and A Comprehensive Multimodal Dataset Towards General Medical AI

### v1

@@ 4.1 Architecture
The GMAI-VL model is a vision-language model built upon the LLaVA architecture [45, 39], incorporating three key components: a large language model (LLM), a vision encoder, and a projector (MLP), as illustrated in Fig. 1(b). These components are designed to work together seamlessly, enabling the model to deliver exceptional performance in medical applications.
We adopt InternLM2.5-7B [63] as our language processing module, which offers outstanding reasoning capabilities. With a context window up to one million tokens, it can handle complex medical tasks and generate coherent, accurate responses. Its support for advanced instruction-following makes it particularly effective in addressing intricate medical queries, thereby enhancing the model’s ability to understand and respond to a wide range of instructions.
For vision processing, GMAI-VL employs a CLIP-based vision encoder [55], which transforms visual inputs into high-dimensional feature representations. CLIP’s strong performance in aligning image and text representations ensures that medical image features are accurately extracted and effectively integrated with linguistic information, significantly enhancing the model’s ability to handle multimodal medical data.
The MLP, as a projector, serves as a bridge between the vision encoder and the LLM, optimizing high-dimensional outputs and further enhancing feature representation. The seamless integration of these components enables GMAI-VL to excel in processing and understanding complex medical multimodal data.
4.2 Optimization Strategy
As illustrated in Fig. 1(c), the training process of the GMAI-VL model is divided into three stages: shadow alignment, deep alignment, instruction tuning, respectively. The detailed hyper-parameter settings can be found in Supp. Mat. (Table 8). To enhance the training of GMAI-VL, we supplement our GMAI-VL-5.5M dataset with additional medical datasets, ultimately forming an 11.7M dataset for training. This supplemented data increases the diversity of training data, exposing the model to a wider range of medical scenarios and visual-language patterns, enhancing its generalization to complex clinical tasks and ensuring robustness in real-world applications. Fig. 5 (in Supp. Mat.) provides a complete distribution of the utilized datasets during the training stage. The detailed data proportions for each training stage are detailed in Table 7 (in Supp. Mat.).
@@ 4.2 Optimization Strategy
In the shallow alignment phase, we utilize a large-scale medical image-text dataset comprising approximately 11.7 million image-text pairs, sourced from a combination of publicly accessible datasets and proprietary in-house data. To achieve shallow alignment, we freeze both the large language model and the vision encoder, optimizing only the projector. With this optimization stage, the model establishes an initial alignment between medical images and their corresponding textual descriptions. All input images are resized to $336\times 336$ pixels, and the training objective is to minimize the cross-entropy loss of the text tokens.
@@
Most vision encoders in multimodal models are pre-trained on natural images, so it is essential to address the domain gap between medical and natural images during the deep alignment stage. To achieve this, we fine-tune both the vision-language projector and the vision encoder, enhancing the alignment between the visual features of medical images and the language model’s feature space.
@@
At this stage, we fine-tune our GMAI-VL model (including the vision encoder, the language model, and the projector components) by instruction tuning to enhance its instruction-following and dialogue capabilities. The multimodal instruction tuning data is primarily derived from the training data in previous stages, with a filtering process to select high-quality and more suitable data for fine-tuning. Specifically, we manually reviewed a subset of each dataset and labeled it as either “high quality” or “low quality.” High-quality datasets were used for tuning, while most low-quality ones were excluded, except for a small portion retained for diversity. This approach allowed efficient filtering without reviewing every sample. Additionally, we incorporate medical text dialogue data to ensure the model’s versatility in handling various dialogue scenarios. Thus, our instruction tuning data comprises approximately ten million samples.
5 Experimental Results
To evaluate our model, we employed several established multimodal medical benchmarks, each targeting specific aspects of medical image understanding and question answering. Below is a brief overview of the benchmarks used in our experiments:
  • Traditional medical VQA benchmarks: Traditional multimodal medical question-answering benchmarks, such as VQA-RAD [38], SLAKE [44], and PMCVQA [80], span various imaging modalities tasks, primarily assessing the model’s ability to extract information from medical images and answer clinical questions. They evaluate the model’s performance in understanding medical imaging and integrating multimodal information.
  • OmniMedVQA: OmniMedVQA [31] provides a rich dataset of paired medical images and text, designed to evaluate the model’s ability to recognize and understand fundamental medical imaging concepts, focusing on cross-modal reasoning and information integration.
  • GMAI-MMBench: GMAI-MMBench [16] focuses on assessing the model’s ability to identify fine-grained objects in complex clinical scenarios, challenging its capacity to handle long-context tasks and accurately recognize and reason over detailed medical features.
  • MMMU Health & Medicine track: The Health & Medicine track of the MMMU [78] benchmark spans a wide range of medical fields, derived from university exams, quizzes, and textbooks. It evaluates the model’s reasoning ability in complex medical scenarios and the specialized knowledge in health and medicine.
5.1 Medical VQA Benchmarks
Table 2: Results on Traditional Medical VQA Benchmarks. The highest performance in each column is highlighted in red, and the second-highest performance is highlighted in blue.
@@ 5.1 Medical VQA Benchmarks
The performance of various VLMs on popular medical VQA benchmark datasets is summarized in Table 2, including VQA-RAD[38], SLAKE[44], and PMC-VQA[80]. Our model, GMAI-VL, demonstrates a strong performance, achieving the highest score on the VQA-RAD[38] dataset with 66.3%, outperforming other models such as HuatuoGPT-Vision-7B. This result highlights GMAI-VL’s superior capability in handling radiological image question-answering tasks. For the PMC-VQA[80] dataset, GMAI-VL achieves 54.3%, and 72.9% on SLAKE[44], demonstrating its capability in handling medical VQA tasks across diverse modalities.
GMAI-VL demonstrates competitive performance across multiple benchmarks, showcasing its versatility in medical image understanding and question-answering.
5.2 OmniMedVQA
Table 3: Comparison of performance between representative LVLMs and GMAI-VL on OmniMedVQA across five different question type. The best performance in each column is highlighted in red, and the second-best performance is highlighted in blue. Abbreviations: MR = Modality Recognition, AI = Anatomy Identification, DD = Disease Diagnosis, LG = Lesion Grading, OBA = Other Biological Attributes.
@@ 5.2 OmniMedVQA
Table 3 summarizes the performance of various large vision-language models (LVLMs), including our proposed GMAI-VL, across five question types: Modality Recognition, Anatomy Identification, Disease Diagnosis, Lesion Grading, and Other Biological Attributes. GMAI-VL demonstrates outstanding accuracy across multiple tasks, achieving 98.64% in Modality Recognition, 92.95% in Anatomy Identification, and 88.71% in Disease Diagnosis. It outperforms both open-source LVLMs and medical-specific models, underscoring its capability to accurately identify anatomical structures and diagnose diseases from visual data. In Lesion Grading, GMAI-VL attained the highest score of 87.21%, and it also delivers a strong performance of 82.95% in Other Biological Attributes, showcasing its versatility across diverse biological contexts. With an average accuracy of 88.48%, the highest among all evaluated models, GMAI-VL excels not only in general medical question-answer tasks but also in complex reasoning requiring domain-specific knowledge, surpassing models like HuatuoGPT-Vision-34B and InternVL2-40B.
These results verify our GMAI-VL is a leading model in multimodal medical image understanding, setting a new benchmark for medical VQA tasks. Its consistent top performance across question types highlights its potential for broader applications in medical question answering.
5.3 GMAI-MMBench
Table 4: Results on the val and test sets of GMAI-MMBench for clinical VQA tasks. The full names of the evaluated tasks can be found in Table 5 in literature [16]. The best model in each category is highlighted in red, while the second-best model is indicated in blue.
@@ 5.3 GMAI-MMBench
Random Guess
Open-Source LVLMs
Proprietary LVLMs
@@
The GMAI-MMBench benchmark is a comprehensive medical multimodal benchmark designed to evaluate models on a range of clinical visual question-answering (VQA) tasks. Table 4 presents the results of various LVLMs, including open-source LVLMs and commercial models, evaluated on the val and test sets across multiple clinical tasks.
Notably, GMAI-VL excels in specific tasks such as abnormality recognition (73.78%), biological variation recognition (63.06%), and clinical disease diagnosis (66.67%). These results demonstrates the model’s strong ability in understanding and interpreting complex clinical images. In comparison to other models, GMAI-VL consistently achieves either the best or second-best performance across most tasks. For instance, it ranks first in 16 out of 20 categories, including key tasks such as AR (Attribute Recognition) and DD (Disease Diagnosis), where it achieved scores of 75.26% and 67.14%, respectively, suggesting GMAI-VL’s strength in understanding medical scenarios.
Overall, GMAI-VL establishes a new benchmark in various clinical VQA tasks, demonstrating its potential as a reliable and versatile tool in medical multimodal applications.
5.4 MMMU Health & Medicine track
The MMMU benchmark, a widely recognized standard for evaluating multimodal models, was used to assess the performance of our proposed GMAI-VL model on the Health & Medicine track. The experimental results in Table 5 show the model’s performance across five key categories: Basic Medical Science (BMS), Clinical Medicine (CM), Diagnostics and Laboratory Medicine (DLM), Pharmacy (P), and Public Health (PH). GMAI-VL performs strongly across multiple categories, achieving top scores in DLM (43.3%), P (50.0%), and PH (53.3%), surpassing competitive models like LLaVA-v1.6 and HuatuoGPT-Vision-7B. These results highlight the model’s proficiency in handling complex tasks requiring diagnostic reasoning, pharmaceutical knowledge, and public health expertise. In BMS, GMAI-VL scores 50.0%, achieve the best performance, demonstrating the model’s the capacity of understanding medical knowledge. In CM, the model achieves 60.0%, remaining competitive with other leading models. These results underscore the model’s ability in processing both clinical and foundational medical information effectively. Overall, GMAI-VL achieves an average score of 51.3% across the Health & Medicine track, which is a top performance among other models, verifying its versatility in specialized medical domains.
Table 5: Performance on the val set for the MMMU Health & Medicine track. This track is divided into five categories: BMS (Basic Medical Science), CM (Clinical Medicine), DLM (Diagnostics and Laboratory Medicine), P (Pharmacy), and PH (Public Health). The best performance in each column is highlighted in red, and the second-best performance is highlighted in blue.
@@ 5.4 MMMU Health & Medicine track
5.5 Ablation Study on GMAI-VL-5.5M Dataset
@@
We further analyzed the effectiveness of the proposed GMAI-VL-5.5M dataset. By removing GMAI-VL-5.5M from the training set, we observed a significant performance drop, as shown in Table 2, Table 3, Table 4 and Table 5. In the VQA-RAD, SLAKE, and PMC-VQA benchmarks, removing GMAI-VL resulted in an average performance improvement of 8.6%. For the OmniMedVQA, GMAI-MMBench, and MMMU Health & Medicine track, the improvements were 8.52%, 6.75%, and 5.3%, respectively. These results demonstrate that the GMAI-VL-5.5M dataset provides highly accurate and reliable medical knowledge, particularly in recognizing and understanding multimodal medical data, which significantly boosts model performance. This underscores the dataset’s diversity, comprehensiveness, and its ability to effectively complement other datasets in complex medical tasks.
@@ 6 Conclusion
In this study, we develop GMAI-VL, a large vision-language model, along with GMAI-VL-5.5M, a comprehensive multimodal medical dataset aimed at advancing general medical AI (GMAI). GMAI-VL-5.5M, which converts hundreds of medical image analysis datasets into high-quality image-text pairs through annotation-guided data generation, enables GMAI-VL to effectively address a wide range of clinical tasks. Experimental results show that GMAI-VL-5.5M significantly enhances GMAI-VL’s performance on diverse clinical tasks, achieving state-of-the-art results across several key benchmarks.
@@
7 Expanded Data Generation Example
In this section, we present the full version of Fig. 2, which provides a detailed example of the data generation process, as shown in Fig.3. This expanded illustration includes the complete prompt and the corresponding detailed responses, enabling a clear comparison between data generation methodologies with and without annotation guidance. The detailed comparison allows for a more intuitive understanding of the distinctions between both methods.
Figure 3: The full version of Fig. 2 in the main text illustrates the complete of data generation pipeline, comparing without-annotation-guided and annotation-guided methods. The annotation-guided approach integrates specific annotation information (e.g., <image, modality, label, department, bbox [optional]>) to generate high-quality, accurate descriptions, while the without-annotation-guided approach often results in lower-quality outputs.
8 Details of GMAI-VL-5.5M
In Table. 6, we provide sub-datasets information of the multimodal dataset GMAI-VL-5.5M we have constructed. Based on the different data forms introduced in the paper, we have categorized the data into five distinct sub-datasets. These include GMAI-MM-Caption-1.7M, GMAI-MM-Instruct-0.9M, GMAI-MM-Percept-1.3M, GMAI-Text-Single-1M, and GMAI-Text-0.6M. Each sub-dataset corresponds to specific components: image caption data, free instruction data, visual perception data, text-only, and conversation. Furthermore, Fig. 4 presents a comprehensive distribution of the data across different modalities, original tasks, departments, and clinical task within this multimodal dataset, highlighting its extensive richness and diversity.
Table 6: Sub-Dataset Details for GMAI-VL-5.5M
@@ 7 Expanded Data Generation Example
9 Training Data Overview
In this section, we provide a comprehensive overview of all datasets utilized for training the GMAI-VL model. The details include the dataset names, their corresponding categories, the amount of data used for training, and the proportion of training data allocated to each dataset during the three phases of model training. Fig. 5 visualizes the distribution of all our training data, showing the proportion of each category and subcategory.
Table. 7 summarizes the datasets employed, highlighting their respective categories and sizes. It is important to note that for certain datasets, we performed data cleaning and bilingual translation. As a result, the dataset sizes reported here may differ from the official numbers.
@@ 9 Training Data Overview
Figure 4: Distribution of GMAI-VL-5.5M across tasks, modalities, departments, and clinical tasks. (a) Original Task Distribution: The dataset includes 2D Classification (50.4%), 3D Segmentation (30.3%), 2D Segmentation (12.7%), and 2D Detection (6.6%). (b) Modality Distribution: In addition to CT (26.8%) and MR (24.7%), X-ray (12.6%), Pathology (11.2%), and less common modalities like Dermoscopy (3.5%), Microscopy (2.4%), and PET (0.2%) are represented. (c) Department Distribution: While Orthopedic Surgery (12.9%) and General Surgery (10.3%) are the top contributors, departments like Endocrinology (1.3%), Infectious Diseases (0.8%), and Urology (0.7%) also provide data. (d) Clinical Task Distribution: Besides Disease Diagnosis (40.4%) and Organ Recognition (16.0%), tasks such as Muscle Recognition (3.3%), Nervous Tissue Recognition (1.5%), and Microorganism Recognition (1.2%) are included. Note: The distribution statistics shown here pertain only to the multimodal components of GMAI-VL-5.5M.
Figure 5: Distribution of all our training data. The inner ring represents major categories, each depicted in a different color. The outer ring corresponds to the subcategories within each major category. The size of each segment is proportional to the amount of data, as indicated in the legend, where the data volume for each subcategory is also provided.

### v2

@@ 4.1 Architecture
The GMAI-VL model is a vision-language model built upon the LLaVA architecture [45, 40], incorporating three key components: a large language model (LLM), a vision encoder, and a projector (MLP), as illustrated in Fig. 1(b).
We use InternLM2.5-7B [62] as our language processing module, which provides exceptional reasoning capabilities. With a context window of up to one million tokens, it can manage complex medical tasks and generate coherent, accurate responses. For vision processing, we adopt a CLIP-based vision encoder [54], which converts visual inputs into high-dimensional feature representations. The MLP projector bridges the vision encoder and language processing module, optimizing high-dimensional outputs and improving feature representation. This framework enables GMAI-VL to effectively process multimodal medical data.
4.2 Training Strategy
As shown in Fig.1(c), the training process of the GMAI-VL model is divided into three stages: shallow alignment, deep alignment, and instruction tuning. To enhance the training of GMAI-VL, we supplement the GMAI-VL-5.5M dataset with additional medical datasets, resulting in a total training dataset of 11.7M samples.
@@ 4.2 Training Strategy
In this phase, we use a large-scale medical image-text dataset of approximately 11.7 million image-text pairs. During this stage, we freeze both the large language model and the vision encoder, optimizing only the projector. This optimization establishes an initial alignment between medical images and their corresponding textual descriptions.
During training, (i) the objective is to minimize the cross-entropy loss of the text tokens; (ii) all images are resized to $336\times 336$ pixels; (iii) the learning rate is set to $1e^{-3}$ with a cosine decay schedule, and AdamW is used as the optimizer; (iv) the total batch size is $32\times 8\times 2$, where 32 refers to the number of GPUs used, eight represents the micro-batch size per GPU, and two is the number of gradient accumulation steps; and (v) a soft packing technique is used to allow each sample to contain multiple sequences, averaging over two sequences per sample.
@@
Most vision encoders in multimodal models are pre-trained on natural images, so addressing the domain gap between medical and natural images is critical during the deep alignment stage. To bridge this gap, we fine-tune both the vision-language projector and the vision encoder, enhancing the alignment between the visual features of medical images and the language model’s feature space. The learning rate is set to $1e^{-4}$, the total batch size is $32\times 4\times 4$, and other settings remain consistent with Stage I.
@@
In this stage, we fine-tune the entire GMAI-VL model—vision encoder, language model, and projector—through instruction tuning to enhance its instruction-following and dialogue capabilities. The multimodal instruction data is primarily derived from the previous stages. Note that this stage discards the low-quality samples that with too short descriptions or extremely lower confidence in the previous stages. Additionally, medical text dialogue data is incorporated to improve the model’s handling of various dialogue scenarios. This results in ten million samples for instruction tuning. During taining, the learning rate is set to $1e^{-5}$, the total batch size is $32\times 4\times 4$, while other parameters, including optimizer, remain unchanged across stages.
Table 4: Results on medical VQA benchmarks. The highest performance in each column is highlighted in red, and the second-highest performance is highlighted in blue.
@@
5 Experimental Results
To evaluate our model, we utilized several established multimodal medical benchmarks, including medical VQA benchmarks (PMCVQA [77], PathVQA [30], VQA-RAD [39], and SLAKE [44]) as well as benchmarks designed for large vision-language models (OmniMedVQA [32], GMAI-MMBench [16], and the MMMU Health & Medicine track [75]). These benchmarks target specific aspects of medical image understanding and question answering. Detailed information about these benchmarks can be found in the supplementary material.
We evaluate model performance using VLMEvalKit [26] with its default settings. To prevent data leakage in GMAI-VL-5.5M, we ensure that the evaluation datasets, OmniMedVQA and GMAI-MMBench, are sourced exclusively from public test sets, while the training data is strictly from public training sets. Additionally, MD5 hashes were computed for each image to verify that there are no duplicates with the benchmark images.
5.1 Comparisons on Medical VQA Datasets
The performance of various VLMs on popular medical VQA benchmark datasets is summarized in Table 4. Our model, GMAI-VL (ours), demonstrates strong performance, achieving the highest score of 66.3% on the VQA-RAD [39] dataset, outperforming models such as HuatuoGPT-Vision-7B. This result underscores GMAI-VL’s superior capability in handling radiological image question-answering tasks. On the PMC-VQA [77] dataset, GMAI-VL achieves 54.3%, and 72.9% on SLAKE [44], highlighting its effectiveness across diverse medical VQA tasks. Overall, GMAI-VL demonstrates competitive performance across multiple benchmarks, showcasing its versatility in medical image understanding and question answering.
5.2 Comparisons on OmniMedVQA
Table 5: Comparison of LVLMs and GMAI-VL on OmniMedVQA across five question types. The best performance in each column is highlighted in red, and the second-best in blue. Abbreviations: MR = Modality Recognition, AI = Anatomy Identification, DD = Disease Diagnosis, LG = Lesion Grading, OBA = Other Biological Attributes.
@@ 5.2 Comparisons on OmniMedVQA
Table 6: Results on the val and test sets of GMAI-MMBench for clinical VQA tasks. Full task names are in Table 5 of [16], and additional model comparisons are in Table 11 (Supplementary Materials). The best and second-best models are marked in red and blue, respectively.
@@
Table 7: Evaluation on the training strategy.
Model MMMU_val OmniMedVQA GMAI_MMBench_val GMAI_MMBench_test
stage1 46.00 81.38 51.49 53.69
stage1+2 46.67 84.64 55.85 58.28
stage1+3 45.33 86.62 59.25 60.70
stage1+2+3 (Prop) 51.30 88.48 61.74 62.43
Table 8: Performance on the val set for the MMMU Health & Medicine track. The best performance is highlighted in red while the second-best performance is highlighted in blue. Note that the results are obtained from the official website.
@@
Figure 2: Example results of our GMAI-VL model. Figure (e) is a failed case.
The OmniMedVQA [32] benchmark integrates 73 traditional medical imaging datasets, all formatted for visual question answering (VQA). Table 5 summarizes the performance of various large vision-language models (LVLMs), including GMAI-VL, across five question types: Modality Recognition, Anatomy Identification, Disease Diagnosis, Lesion Grading, and Other Biological Attributes.
From the results, we can see that (i) GMAI-VL excels in all tasks, achieving 98.64% in Modality Recognition, 92.95% in Anatomy Identification, and 88.71% in Disease Diagnosis; (ii) it outperforms both open-source and medical-specific models, demonstrating strong abilities in identifying anatomical structures and diagnosing diseases; (iii) in Lesion Grading, it achieves the highest score of 87.21%, and scores 82.95% in Other Biological Attributes, showcasing its versatility; and (iv) with an average accuracy of 88.48%, GMAI-VL surpasses models like HuatuoGPT-Vision-34B and InternVL2-40B, establishing itself as a leading model in multimodal medical image understanding and setting a new benchmark for medical VQA tasks.
5.3 Comparisons on GMAI-MMBench
The GMAI-MMBench benchmark is a comprehensive evaluation suite for medical multimodal models, focusing on clinical visual question-answering (VQA) tasks. Table 6 presents the performance of various models on the val and test sets across a range of clinical tasks. Notably, (i) GMAI-VL excels in tasks such as abnormality recognition (73.78%), biological variation recognition (63.06%), and clinical disease diagnosis (66.67%), demonstrating its strong ability to understand and interpret complex clinical images. (ii) Compared to other models, GMAI-VL consistently ranks first or second across most tasks, securing the top position in 16 out of 20 categories. (iii) Key tasks such as Attribute Recognition (AR) and Disease Diagnosis (DD) yield scores of 75.26% and 67.14%, respectively, highlighting GMAI-VL’s strength in medical scenario understanding. (iv) Overall, GMAI-VL sets a new benchmark in various clinical VQA tasks.
5.4 Comparisons on MMMU Health & Medicine
We further assess the performance of our GMAI-VL on the Health & Medicine track of MMMU benchmark, which is a widely recognized standard for evaluating multimodal models. The experimental results in Table 8 show the model’s performance across five key categories: Basic Medical Science (BMS), Clinical Medicine (CM), Diagnostics and Laboratory Medicine (DLM), Pharmacy (P), and Public Health (PH).
The results show that (i) GMAI-VL performs strongly across multiple categories, achieving top scores in DLM (43.3%), P (50.0%), and PH (53.3%), surpassing competitive models such as LLaVA-v1.6 and HuatuoGPT-Vision-7B. These results highlight the model’s proficiency in handling complex tasks that require diagnostic reasoning, pharmaceutical knowledge, and public health expertise. (ii) In BMS, GMAI-VL scores 50.0%, achieving the best performance and demonstrating its ability to understand medical knowledge. (iii) In CM, the model scores 60.0%, remaining competitive with other leading models. These results underscore the model’s effectiveness in processing both clinical and foundational medical information. (iv) Overall, GMAI-VL achieves an average score of 51.3% across the Health & Medicine track, ranking among the top models and confirming its versatility in specialized medical domains.
5.5 Ablation Study of the GMAI-VL-5.5M Dataset
In this section, we evaluate the effectiveness of the proposed GMAI-VL-5.5M dataset. We report the results of GMAI-VL (w/o our data) in Tables 4, 5, 6, and 8. This model is trained using data excluding the GMAI-VL-5.5M dataset (see Sec.4.2), highlighting that our dataset effectively enhances the performance of large vision-language models.
@@ 5.4 Comparisons on MMMU Health & Medicine
The results demonstrate that the GMAI-VL-5.5M dataset provides highly accurate and reliable medical knowledge, especially in recognizing and understanding multimodal medical data. This significantly improves model performance, showcasing the dataset’s diversity, comprehensiveness, and its ability to complement other datasets in complex medical tasks.
5.6 Ablation Study of Training Strategy
Table 7 presents results at different training stages. (i) The proposed stage1+2+3 (Prop) training strategy significantly enhances model performance across multiple benchmarks, outperforming other models and training strategies in all key metrics. (ii) By progressively incorporating stages 1, 2, and 3, the model shows consistent improvements on various datasets. This highlights the effectiveness of the three-stage training pipeline in improving the model’s ability to handle complex multimodal medical tasks, confirming that a more comprehensive training approach yields superior results.
5.7 Case Study
Fig. 2 shows example results of our GMAI-VL across various medical imaging and diagnostic tasks. The model accurately interprets chest X-rays, detects lesions in fundus images, and identifies COVID-19 features in chest CT scans, among others. These results highlight the model’s versatility and potential in assisting clinical diagnostics. However, as illustrated in the failure case in Fig. 2 (e), GMAI-VL encounters difficulties in detecting subtle differences, such as distinguishing between mild and moderate severity.
@@ 6 Conclusion
In this paper, we build GMAI-VL, a large vision-language model, and GMAI-VL-5.5M, a comprehensive multimodal medical dataset designed to advance general medical AI. GMAI-VL-5.5M converts hundreds of medical image analysis datasets into high-quality image-text pairs through annotation-guided data generation, enabling GMAI-VL to tackle a wide range of clinical tasks effectively. Experimental results show that GMAI-VL-5.5M significantly enhances GMAI-VL’s performance, achieving state-of-the-art results across multiple key benchmark datasets. Future work will focus on expanding the dataset with more diverse and challenging medical scenarios, further improving the model’s ability to generalize across different clinical environments and applications.
{CJK}
UTF8gbsn
@@
GMAI-VL & GMAI-VL-5.5M: A Large Vision-Language Model and a Comprehensive Multimodal Dataset Towards General Medical AI
There are four parts in this supplementary material.
Part 1: Details of GMAI-VL-5.5M
Part 2: Details of All Training Data for GMAI-VL
Part 3: Results on the Validation and Test Sets of GMAI-MMBench for Clinical VQA Tasks
Part 4: Quality Evaluation
Part 1 - Details of GMAI-VL-5.5M
Table 9: Details of Sub-Datasets in GMAI-VL-5.5M
@@
Figure 3: Distribution of GMAI-VL-5.5M across various modalities, departments, and tasks.
Table 9 provides information on the sub-datasets of the multimodal dataset GMAI-VL-5.5M that we have constructed. Based on the different data formats discussed in the paper, we have categorized the data into five distinct sub-datasets: GMAI-MM-Caption-1.7M, GMAI-MM-Instruct-0.9M, GMAI-MM-Percept-1.3M, GMAI-Text-Single-1M, and GMAI-Text-0.6M. Each sub-dataset corresponds to specific components, including image caption data, free instruction data, visual perception data, text-only data, and conversation data. Additionally, Fig. 3 presents a comprehensive distribution of the data across different modalities, original tasks, departments, and clinical tasks within the dataset, highlighting its richness and diversity.
Part 2 - Details of All Training Data for GMAI-VL
Figure 4: Distribution of the training dataset. The inner ring represents major categories, each shown in a distinct color, while the outer ring depicts the corresponding subcategories. Segment sizes are proportional to data volume, as indicated in the legend, which also provides the data volume for each subcategory.